# Sonic4D: Spatial Audio Generation for Immersive 4D Scene Exploration

### v1

@@ Abstract
Recent advancements in 4D generation have demonstrated its remarkable capability in synthesizing photorealistic renderings of dynamic 3D scenes. However, despite achieving impressive visual performance, almost all existing methods overlook the generation of spatial audio aligned with the corresponding 4D scenes, posing a significant limitation to truly immersive audiovisual experiences. To mitigate this issue, we propose Sonic4D, a novel framework that enables spatial audio generation for immersive exploration of 4D scenes. Specifically, our method is composed of three stages: 1) To capture both the dynamic visual content and raw auditory information from a monocular video, we first employ pre-trained expert models to generate the 4D scene and its corresponding monaural audio. 2) Subsequently, to transform the monaural audio into spatial audio, we localize and track the sound sources within the 4D scene, where their 3D spatial coordinates at different timestamps are estimated via a pixel-level visual grounding strategy. 3) Based on the estimated sound source locations, we further synthesize plausible spatial audio that varies across different viewpoints and timestamps using physics-based simulation. Extensive experiments have demonstrated that our proposed method generates realistic spatial audio consistent with the synthesized 4D scene in a training-free manner, significantly enhancing the immersive experience for users. Generated audio and video examples are available at https://x-drunker.github.io/Sonic4D-project-page.
1 Introduction
Benefiting from large-scale data available soomro2012ucf101 ; schuhmann2021laion ; yu2023celebv and recent advancement in generative models ho2020denoising ; song2020denoising , 4D generation (i.e., dynamic 3D scene generation) jiang2024consistentd ; ren2023dreamgaussian4d ; zeng2024stag4d ; zhu2025ar4d has emerged as a promising direction due to its powerful capability in modeling complex spatiotemporal dynamics of real-world scenes. By enabling spatiotemporally consistent renderings from arbitrary camera viewpoints, 4D generation facilitates various downstream applications such as AR/VR li20244k4dgen ; fritsch20173d , robotics khalid20224d ; hann20204d , and autonomous driving wang2024occsora ; min2024driveworld .
However, while existing 4D generation methods hann20204d ; liu2025free4d ; zhang20244diffusion ; zeng2024trans4d achieve impressive visual results, they typically neglect the generation of spatial audio consistent with the corresponding 4D scene (i.e., audio that varies with the listener’s viewpoint and follows physical acoustic principles), limiting the overall immersive experience to a large extent.
@@ 1 Introduction
  • We propose a three-stage framework to achieve spatial audio generation: 1) Dynamic Scene and Monaural Audio Generation to extract semantically aligned visual and audio priors from a single video; 2) 3D Sound-Source Localization and Tracking to recover sound source’s 3D trajectory for precise acoustic simulation; 3) Physics-Driven Spatial Audio Synthesis to render dynamic, viewpoint-adaptive binaural audio via physics-based room impulse response simulation.
@@
2 Related Works
2.1 4D Generation
Early advances in 4D generation have built upon techniques such as Score Distillation Sampling (SDS) poole2022sds , which leverages generative priors of pre-trained diffusion models to optimize dynamic scene representations. For example, MAV3D singer2023text4d proposed a multi-stage framework that sequentially optimizes static geometry using text-to-image diffusion, followed by motion synthesis using text-to-video models. This paradigm laid the groundwork for methods such as Dream-in-4D zheng2024unified , which further integrated motion priors from pretrained video diffusion models into dynamic NeRF-based representations pumarola2021d ; yan2023nerf . To improve rendering efficiency and motion fidelity, several methods have evolved from volumetric models to more expressive and compact formulations, such as AYG ling2024alignyourgaussians and Trans4D zeng2024trans4d , which employ deformation networks to animate static Gaussians across time. Concurrently, models like Consistent4D jiang2024consistentd and 4Diffusion zhang20244diffusion focused on enhancing spatiotemporal consistency, utilizing multiview interpolation, temporal alignment modules, or synchronized training strategies. Works such as Diffusion4D liang2024diffusion4d , PLA4D miao2024pla4d , and 4Dynamic yuan20244dynamic further emphasized controllability and efficiency by leveraging pixel-aligned supervision, video-based guidance, or hybrid representations combining mesh and Gaussian structures. Single-image driven methods like DreamGaussian4D ren2023dreamgaussian4d , Human4DiT shao2024human4dit , and Animate124 zhao2023animate124 extended monocular inputs into coherent 4D outputs using hierarchical architectures or diffusion-based motion refinement. Recent studies such as Free4D liu2025free4d and MVTokenFlow huang2025mvtokenflow further improve spatial-temporal coherence and visual quality without tuning large generative models.
@@ 2 Related Works
2.2 Spatial Audio Generation
@@
Early spatial audio generation methods  gao20192 ; xu2021visually ; leng2022binauralgrad predominantly took monaural inputs and “binauralized” them via neural networks conditioned on visual cues. In recent years, the advent of powerful generative architectures has spurred end-to-end models that accept text prompts sun2024both , images sun2024both ; dagli2024see , spatial parameters heydari2025immersediffusion ; kushwaha2025diff , silent videos kim2025visage , or full 360° panoramas liu2025omniaudio to jointly learn semantic and spatial audio features. For example, SpatialSonic sun2024both leverages spatial-aware encoders and azimuth state matrices within a latent diffusion framework to provide fine-grained multimodal spatial guidance for stereo audio generation. ViSAGe kim2025visage fuses CLIP visual embeddings with autoregressive neural audio codec modeling to generate coherent FOA directly from silent video frames. Recently, OmniAudio liu2025omniaudio addressed the 360° spatial audio gap by using a dual-branch architecture to fuse panoramic and FoV streams, enabling high-fidelity FOA generation from full spherical video content.
@@
Although these methods benefit from the inherent capacity of generative models to capture complex multimodal distributions and yield impressive qualitative gains, they face two critical challenges. First, the data scarcity and domain gaps characteristic of end-to-end learning frameworks still constrain their ability to faithfully simulate spatial audio, resulting in suboptimal spatial rendering fidelity. Second, even when utilizing 360° video inputs, the synthesized sound field remains tied to the camera’s fixed pose and cannot account for dynamic listener motion or free-viewpoint changes within the scene.
Some alternative approaches chen2023novel ; liang2023av ; chen2024av tackle spatial audio by jointly reconstructing novel views and associated audio given ground-truth recordings at specific viewpoints; these methods, however, are restricted to static scenes and focus on view-specific audio re-rendering. In contrast, our work decouples semantic audio synthesis from spatial rendering and employs room impulse response (RIR) convolution to robustly simulate the spatial acoustics perceived by a moving listener in dynamic environments.
2.3 Sound Source Localization
@@
Sound source localization aims to find the location of sound sources in an image or video frame. Initial studies senocak2018learning ; chen2021localizing ; senocak2023sound ; um2023audio addressed this problem by leveraging methods like cross-modal attention and contrastive learning to establish effective alignment between audio and visual modalities, mainly focusing on single sound source localization. Some works have taken this further to achieve multi-sound source localization hu2020discriminative ; mo2023audio ; kim2024learning , enabling the simultaneous localization of multiple sound sources from mixed audio and visual inputs. However, these approaches focus on exploring audio–visual alignments and cannot localize sound sources without ground-truth audio. Moreover, due to the complexity of three- and even four-dimensional scenarios, few works he2024sound3dvdet has addressed sound source detection or localization in 3D space.
Recently, the rapid advancement of Multimodal Large Language Models zhao2023bubogpt ; munasinghe2024videoglamm ; li2024groundinggpt (MLLMs) has made it possible to perform audio-visual grounding with strong generalization capabilities. In this work, we leverage the powerful prior knowledge of pretrained MLLMs to directly perform pixel-level visual grounding on the video input for sound source localization, followed by back-projection into 3D space to achieve accurate 3D sound source positioning and further spatial audio simulation.
3 Methods
Figure 1: Pipeline of Sonic4D. Our method is composed of three stages: 1) Dynamic Scene and Monaural Audio Generation: extracting semantically aligned visual scenes and audio priors from monocular videos (Sec. 3.1); 2) 3D Sound-Source Localization and Tracking: estimating the sound source’s trajectory in 3D space for physically accurate sound propagation modeling (Sec. 3.2); 3) Physics-Driven Spatial Audio Synthesis: leveraging a physics-based room impulse response simulation to realize spatial audio simulation. (Sec. 3.3)
@@ 3 Methods
3.1 Stage I: Dynamic Scene and Monaural Audio Generation
To obtain a novel-view video $\bm{V}^{r}$ while simultaneously capturing dynamic spatial priors for spatial audio simulation, a model that can produce high-fidelity, free-viewpoint renderings of dynamic scenes with strong spatiotemporal consistency is required. Motivated by recent progress in 4D generation, we adopt TrajectoryCrafter yu2025trajectorycrafter , a pretrained video generative model capable of synthesizing high-fidelity videos along arbitrary camera trajectories from a single-view input video, as our 4D content generator.
Specifically, we first apply a monocular depth estimator to estimate the depth maps $\bm{D}^{s}=\{\bm{D}^{s}_{i}\}_{i=1}^{n}\in\mathbb{R}^{n\times h\times w}$ of the input video $\bm{V}^{s}$. Subsequently, we lift $\bm{V}^{s}$ into a set of dynamic point clouds $\bm{P}=\{\bm{P}_{i}\}_{i=1}^{n}$ using inverse perspective projection $\bm{\Phi}^{-1}$, which is formulated as follows:
@@ 3.1 Stage I: Dynamic Scene and Monaural Audio Generation
Next, we prepare the raw monaural audio signal required for subsequent physics-based spatial audio simulation. To this end, we adopt MMAudio cheng2024taming , a transformer-based video-to-audio synthesis model, to generate monaural audio $\bm{A}^{m}\in\mathbb{R}^{1\times T}$ that is temporally and semantically aligned with the input video $\bm{V}^{s}$. By decomposing the task of immersive 4D scene exploration into two sub-tasks—i.e., 4D scene generation and spatial audio synthesis—and leveraging pretrained expert models to extract the required visual and acoustic information, we can achieve greater flexibility and continually improve performance as upstream models evolve.
3.2 Stage II: 3D Sound Source Localization and Tracking
Figure 2: Illustration of Stage II: 3D Sound Source Localization and Tracking. We feed the video input frame-wisely into GroundingGPT li2024groundinggpt to obtain the sound source pixel coordinates. Then, we compute the average 3D coordinates of an $r\times r$ patch centered around the estimated 2D location via back-projection, and treat this as the 3D position of the sound source. Finally, we apply the DBSCAN ester1996density algorithm to filter and interpolate the sound-source trajectory, yielding a smooth and reliable final trajectory.
With the dynamic point cloud $\bm{P}$ and its associated monaural audio $\bm{A}^{m}$, we proceed to localize and track the sound source in 3D space to facilitate physics-based spatial audio generation. However, directly tracking the sound source in a 4D environment remains highly challenging, particularly in highly dynamic scenes with complex geometry. To address this, we first localize the sound source in the 2D pixel space and then back-project the estimated 2D coordinates into 3D space, thereby modeling the sound source trajectory within the generated dynamic scene.
To achieve this goal, existing 2D sound-source localization methods um2023audio ; senocak2023sound ; park2024can typically require ground-truth audio inputs and are often tailored to specific scenarios and input sizes. This dependence on “true” audio runs counter to Sonic4D’s very generation-centric design. In contrast, given the rich commonsense knowledge and reasoning abilities of multimodal large language models (MLLMs), we propose to leverage GroundingGPT li2024groundinggpt , a highly generalizable MLLM focused on grounding tasks and trained on a variety of audio–visual datasets for sound source localization, as our 2D sound source localization component.
Concretely, as shown in Fig. 1, we craft a textual prompt:
In this image, there is a possible sound source. Please find it and estimate its position. The only value you need to return is the position of the sound source in the format [x0, y0, x1, y1]. Do not include any text other than that.
and feed, frame by frame, the source video $\bm{V}^{s}$ together with the prompt into GroundingGPT li2024groundinggpt . Assuming the grounding result for the $i$-th frame $\bm{I}_{i}^{s}$ is represented by a bounding box $[x_{0},y_{0},x_{1},y_{1}]$, we define the sound source pixel coordinates as the center of this bounding box, i.e.:
@@ 3.2 Stage II: 3D Sound Source Localization and Tracking
Here, $\pi^{-1}$ denotes the back-projection operator that maps a pixel coordinate and its depth into a 3D point. Considering the sporadic errors introduced by monocular depth estimation at individual pixels, we compute the 3D coordinates of an $r\times r$ pixel patch surrounding $(u_{i},v_{i})$ and average them to obtain the sound-source’s 3D location as shown in Figure 2:
@@
where $\mathcal{P}^{r}(u_{i},v_{i})$ denotes the set of all pixels $(u,v)$ such that $\lvert u-u_{i}\rvert\leq(r-1)/2$ and $\lvert v-v_{i}\rvert\leq(r-1)/2$. Nonetheless, due to uncertainty in the bounding-box predictions, the back-projected trajectory $\{\bar{\mathbf{X}}_{i}\}$ may still exhibit significant fluctuations. To mitigate this, we apply DBSCAN ester1996density (Density-Based Spatial Clustering of Applications with Noise), a clustering-based outlier detection algorithm that groups dense regions while marking sparse points as noise, to filter out spurious estimates. Points marked as noise are then temporally filled in via linear interpolation between the nearest non–noise neighbours to produce a smooth trajectory $\textbf{Traj}_{src}=\{\bm{X}_{i}\}_{i=1}^{n}$.
In addition to the sound source’s trajectory, the receiver trajectory is also required for acoustic simulation. We use the user-specified camera trajectory as the receiver trajectory. For each frame $i$, the camera pose is defined as
@@
3.3 Stage III: Physics-Driven Spatial Audio Synthesis
Given the sound source’s trajectory $\textbf{Traj}_{src}$, the receiver trajectory $\textbf{Traj}_{rsv}$, and the monaural audio signal $\bm{A}^{m}$, we perform a physics-based spatial audio simulation.
A central element of our simulation is the computation of Room Impulse Responses (RIRs) via the Image Source Method (ISM) allen1979image . Building on the Image-Source Method for static RIRs, we apply gpuRIR diaz2021gpurir to simulate dynamic sources and receivers by segmenting the trajectories into short intervals over which they are assumed stationary. Concretely, let the total signal length be $N$ samples and split it into $M$ segments at sample indices
@@ 3.3 Stage III: Physics-Driven Spatial Audio Synthesis
$h_{i,L}(\tau)=\mathrm{ISM}\bigl{(}\mathbf{x}^{s}_{i},\;\mathbf{x}^{r}_{i,L}\bigr{)},\quad h_{i,R}(\tau)=\mathrm{ISM}\bigl{(}\mathbf{x}^{s}_{i},\;\mathbf{x}^{r}_{i,R}\bigr{)}.$ (7)
@@
Then the left and right binaural signals are obtained by block-wise convolution:
$y_{L}[n]=\sum_{i=0}^{M-1}\bigl{(}b_{i}*h_{i,L}\bigr{)}[\,n-n_{i}\,],\qquad y_{R}[n]=\sum_{i=0}^{M-1}\bigl{(}b_{i}*h_{i,R}\bigr{)}[\,n-n_{i}\,],$ (9)
where $h_{i,L}$ and $h_{i,R}$ are the left- and right-ear impulse responses in $h_{i}$. Putting it all together, the final binaural signal is
@@
To ensure the output audio is in a valid range for playback, we normalize the signal to the range $[-1,1]$:
@@
By decomposing the task into three modular stages, our framework maximizes flexibility and extensibility. Each stage benefits from pretraining or physical priors, allowing Sonic4D to scale with future advances in generative modeling and spatial simulation. Moreover, this decoupled design enables plug-and-play improvements, making Sonic4D suitable for a broad range of content creation and immersive applications.
4 Experiments
4.1 Experimental Settings
Since this is the first attempt to perform spatial audio generation for immersive 4D scene exploration, existing evaluation benchmarks are not directly applicable. Therefore, we conducted a user preference study to assess the perceptual quality of spatial audio generated by our proposed and other methods. We choose MMAudio cheng2024taming as the baseline method, whose generated audio is monaural and contains no spatialization.
@@ 4.1 Experimental Settings
To determine whether our generated spatial audio conveys a sense of space, we paired the same viewpoint-specific video with either the spatialized stereo audio or the non-spatialized mono audio for comparison. Participants were first asked, for each video pair (identical visuals but different audio), to choose which audio track felt more spatial. Next, they rated both audio tracks on two criteria from 1 to 5:
@@
  • Spatial Localization Accuracy: Measures how faithfully the interaural level and time-difference cues recreate the true 3D source position. Participants rate from 1 (“no sense of source direction”) to 5 (“precise and stable perception of source location”).
@@
  • Audio–Visual Spatial Congruence: Assesses how consistently the evolving audio cues (e.g. changes in loudness or binaural disparity) track the on-screen motion of the subject or camera. Participants rate from 1 (“audio movement conflicts with visuals”) to 5 (“audio shifts perfectly match visual motion”).

### v3

# arXiv:2506.15759v3[cs.SD] 27 Feb 2026

@@ Abstract
Recent advancements in 4D generation have demonstrated its remarkable capability in synthesizing photorealistic renderings of dynamic 3D scenes. However, despite achieving impressive visual performance, almost all existing methods overlook the generation of spatial audio aligned with the corresponding 4D scenes, posing a significant limitation to truly immersive audiovisual experiences. To mitigate this issue, we propose Sonic4D, a novel framework that enables spatial audio generation for immersive exploration of 4D scenes. Specifically, our method is composed of three stages: 1) To capture both the dynamic visual content and raw auditory information from a monocular video, we first employ pre-trained expert models to generate the 4D scene and its corresponding monaural audio. 2) Subsequently, to transform the monaural audio into spatial audio, we localize and track the sound sources within the 4D scene, where their 3D spatial coordinates at different timestamps are estimated via a pixel-level visual grounding strategy. 3) Based on the estimated sound source locations, we further synthesize plausible spatial audio that varies across different viewpoints and timestamps using physics-based simulation. Extensive experiments have demonstrated that our proposed method generates realistic spatial audio consistent with the synthesized 4D scene in a training-free manner, significantly enhancing the immersive experience for users.
Introduction
Benefiting from large-scale data available (Soomro et al. 2012; Schuhmann et al. 2021; Yu et al. 2023) and recent advancement in generative models (Ho et al. 2020; Song et al. 2020), 4D generation (i.e., dynamic 3D scene generation) (Jiang et al. 2024; Ren et al. 2023; Zeng et al. 2024b; Zhu et al. 2025) has emerged as a promising direction due to its powerful capability in modeling complex spatiotemporal dynamics of real-world scenes. By enabling spatiotemporally consistent renderings from arbitrary camera viewpoints, 4D generation facilitates various downstream applications such as AR/VR (Li et al. 2024a; Fritsch and Klein 2017), robotics (Khalid et al. 2022; Hann et al. 2020), and autonomous driving (Wang et al. 2024; Min et al. 2024).
However, while existing 4D generation methods (Hann et al. 2020; Liu et al. 2025b; Zhang et al. 2024; Zeng et al. 2024a) achieve impressive visual results, they typically neglect the generation of spatial audio consistent with the corresponding 4D scene (i.e., audio that varies with the listener’s viewpoint and follows physical acoustic principles), limiting the overall immersive experience to a large extent.
@@ Introduction
  • We propose a three-stage framework to achieve spatial audio generation: 1) Dynamic Scene and Monaural Audio Generation to extract semantically aligned visual and audio priors from a single video; 2) 3D Sound-Source Localization and Tracking to recover sound source’s 3D trajectory for precise acoustic simulation; 3) Physics-Driven Spatial Audio Synthesis to render dynamic, viewpoint‑adaptive binaural audio via physics-based room impulse response simulation.
@@
Figure 1: Pipeline of Sonic4D. Our method is composed of three stages: 1) Dynamic Scene and Monaural Audio Generation: extracting semantically aligned visual scenes and audio priors from monocular videos; 2) 3D Sound-Source Localization and Tracking: estimating the sound source’s trajectory in 3D space for physically accurate sound propagation modeling; 3) Physics-Driven Spatial Audio Synthesis: leveraging a physics-based room impulse response simulation to realize spatial audio simulation.
Related Work
4D Generation
Early advances in 4D generation have built upon techniques such as Score Distillation Sampling (SDS) (Poole et al. 2022), which leverages generative priors of pre-trained diffusion models to optimize dynamic scene representations. This paradigm laid the groundwork for methods such as Dream-in-4D (Zheng et al. 2024), which further integrated motion priors from pretrained video diffusion models into dynamic NeRF-based representations (Pumarola et al. 2021; Yan et al. 2023). Concurrently, models like Consistent4D (Jiang et al. 2024) and 4Diffusion (Zhang et al. 2024) focused on enhancing spatiotemporal consistency, utilizing multiview interpolation, temporal alignment modules, or synchronized training strategies. Works such as Diffusion4D (Liang et al. 2024), PLA4D (Miao et al. 2024), and 4Dynamic (Yuan et al. 2024) further emphasized controllability and efficiency by leveraging pixel-aligned supervision, video-based guidance, or hybrid representations combining mesh and Gaussian structures. Recent studies such as Free4D (Liu et al. 2025b) and MVTokenFlow (Huang et al. 2025) further improve spatial-temporal coherence and visual quality without tuning large generative models.
@@
Spatial Audio Generation
@@
Early spatial audio generation methods  (Gao and Grauman 2019; Xu et al. 2021; Leng et al. 2022) predominantly took monaural inputs and “binauralized” them via neural networks conditioned on visual cues. In recent years, the advent of powerful generative architectures has spurred end‑to‑end models that accept text prompts (Sun et al. 2024), images (Sun et al. 2024; Dagli et al. 2024), spatial parameters (Heydari et al. 2025; Kushwaha et al. 2025), silent videos (Kim et al. 2025), or full 360° panoramas (Liu et al. 2025a) to jointly learn semantic and spatial audio features. For example, SpatialSonic (Sun et al. 2024) leverages spatial‑aware encoders and azimuth state matrices within a latent diffusion framework to provide fine‑grained multimodal spatial guidance for stereo audio generation. ViSAGe (Kim et al. 2025) fuses CLIP visual embeddings with autoregressive neural audio codec modeling to generate coherent FOA directly from silent video frames. Recently, OmniAudio (Liu et al. 2025a) addressed the 360° spatial audio gap by using a dual‑branch architecture to fuse panoramic and FoV streams, enabling high‑fidelity FOA generation from full spherical video content.
@@
Although these generative methods achieve impressive multimodal results, they still face two key limitations. First, data scarcity and domain gaps in end-to-end learning restrict the fidelity of spatial audio rendering. Second, even with 360° video inputs, the synthesized sound field is bound to a fixed camera pose and cannot support dynamic listener motion or free-viewpoint changes. In contrast, our method decouples semantic audio synthesis from spatial rendering and leverages RIR-based convolution to physically simulate spatial acoustics for a moving listener in dynamic scenes.
Sound Source Localization
@@
Sound source localization aims to find the location of sound sources in an image or video frame. Initial studies (Senocak et al. 2018; Chen et al. 2021; Senocak et al. 2023; Um et al. 2023) addressed this problem by leveraging methods like cross-modal attention and contrastive learning to establish effective alignment between audio and visual modalities, mainly focusing on single sound source localization. Some works have taken this further to achieve multi-sound source localization (Mo and Tian 2023; Kim et al. 2024), enabling the simultaneous localization of multiple sound sources from mixed audio and visual inputs. However, these approaches focus on exploring audio–visual alignments and cannot localize sound sources without ground-truth audio. Recently, the rapid advancement of Multimodal Large Language Models (Munasinghe et al. 2024; Li et al. 2024b) (MLLMs) has made it possible to perform audio-visual grounding with strong generalization capabilities. In this work, we leverage the powerful prior knowledge of pretrained MLLMs to directly perform pixel-level visual grounding on the video input for sound source localization, followed by back-projection into 3D space to achieve accurate 3D sound source positioning and further spatial audio simulation.
Figure 2: Illustration of Stage II: We localize the sound source in each frame using GroundingGPT (Li et al. 2024b), back-project the 2D grounding results to 3D via depth, and apply DBSCAN (Ester et al. 1996) to obtain a smooth trajectory.
Methods
@@
Stage I: Dynamic Scene and Monaural Audio Generation
To obtain a novel-view video $\bm{V}^{r}$ while simultaneously capturing dynamic spatial priors for spatial audio simulation, a model that can produce high-fidelity, free-viewpoint renderings of dynamic scenes with strong spatiotemporal consistency is required. Motivated by recent progress in 4D generation, we adopt TrajectoryCrafter (YU et al. 2025), a pretrained video generative model capable of synthesizing high-fidelity videos along arbitrary camera trajectories from a single-view input video, as our 4D content generator.
Specifically, we first apply a monocular depth estimator (Hu et al. 2025) to estimate the depth maps $\bm{D}^{s}=\{\bm{D}^{s}_{i}\}_{i=1}^{n}\in\mathbb{R}^{n\times h\times w}$ of the input video $\bm{V}^{s}$. Subsequently, we lift $\bm{V}^{s}$ into a set of dynamic point clouds $\bm{P}=\{\bm{P}_{i}\}_{i=1}^{n}$ using inverse perspective projection $\bm{\Phi}^{-1}$, which is formulated as follows:
@@
Next, we prepare the raw monaural audio signal required for subsequent physics-based spatial audio simulation. To this end, we adopt MMAudio (Cheng et al. 2024), a transformer-based video-to-audio synthesis model, to generate monaural audio $\bm{A}^{m}\in\mathbb{R}^{1\times T}$ that is temporally and semantically aligned with the input video $\bm{V}^{s}$. By decomposing the task of immersive 4D scene exploration into two sub-tasks—i.e., 4D scene generation and spatial audio synthesis—and leveraging pretrained expert models to extract the required visual and acoustic information, we can achieve greater flexibility and continually improve performance as upstream models evolve.
Stage II: 3D Sound Source Localization and Tracking
With the dynamic point cloud $\bm{P}$ and its associated monaural audio $\bm{A}^{m}$, we proceed to localize and track the sound source in 3D space to facilitate physics-based spatial audio generation. However, directly tracking the sound source in a 4D environment remains highly challenging. Therefore, we first localize the sound source in the 2D pixel space and then back-project the estimated 2D coordinates into 3D space, thereby modeling the sound source trajectory within the generated dynamic scene.
To achieve this goal, existing 2D sound‑source localization methods (Um et al. 2023; Senocak et al. 2023; Park et al. 2024) typically require ground‑truth audio inputs and are often tailored to specific scenarios and input sizes. This dependence on “true” audio runs counter to Sonic4D’s very generation‑centric design. In contrast, given the rich commonsense knowledge and reasoning abilities of multimodal large language models (MLLMs), we propose to leverage GroundingGPT (Li et al. 2024b), a highly generalizable MLLM focused on grounding tasks and trained on a variety of audio–visual datasets for sound source localization, as our 2D sound source localization component.
Concretely, as shown in Fig. 1, we craft a textual prompt and feed the source video $\bm{V}^{s}$ frame by frame together with the prompt into GroundingGPT (Li et al. 2024b). Assuming the grounding result for the ii‑th frame $\bm{I}_{i}^{s}$ is represented by a bounding box $x_{0},y_{0},x_{1},y_{1}]$, we define the sound source pixel coordinates as the center of this bounding box, i.e.:
@@
Here, $\pi^{-1}$ denotes the back‑projection operator that maps a pixel coordinate and its depth into a 3D point. Considering the sporadic errors introduced by monocular depth estimation at individual pixels, we compute the 3D coordinates of an $r\times r$ pixel patch surrounding $(u_{i},v_{i})$ and average them to obtain the sound‐source’s 3D location as shown in Figure 2:
@@
where $\mathcal{P}^{r}(u_{i},v_{i})$ denotes the set of all pixels $(u,v)$ such that $\lvert u-u_{i}\rvert\leq(r-1)/2$ and $\lvert v-v_{i}\rvert\leq(r-1)/2$. Nonetheless, due to uncertainty in the bounding‐box predictions, the back‐projected trajectory $\{\bar{\mathbf{X}}_{i}\}$ may still exhibit significant fluctuations. To mitigate this, we apply DBSCAN (Ester et al. 1996) (Density‐Based Spatial Clustering of Applications with Noise), a clustering‐based outlier detection algorithm that groups dense regions while marking sparse points as noise, to filter out spurious estimates. Points marked as noise are then temporally filled in via linear interpolation between the nearest non–noise neighbours to produce a smooth trajectory $\textbf{Traj}_{src}=\{\bm{X}_{i}\}_{i=1}^{n}$.
In addition to the sound source’s trajectory, the receiver trajectory is also required for acoustic simulation. We use the user‑specified camera trajectory as the receiver trajectory. For each frame $i$, the camera pose is defined as
@@
Stage III: Physics-Driven Spatial Audio Synthesis
Given the sound source’s trajectory $\textbf{Traj}_{src}$, the receiver trajectory $\textbf{Traj}_{rsv}$, and the monaural audio signal $\bm{A}^{m}$, we perform a physics‑based spatial audio simulation.
A central element of our simulation is the computation of Room Impulse Responses (RIRs) via the Image Source Method (ISM) (Allen and Berkley 1979). Building on the Image‐Source Method for static RIRs, we apply gpuRIR (Diaz-Guerra et al. 2021) to simulate dynamic sources and receivers by segmenting the trajectories into short intervals over which they are assumed stationary. Concretely, let the total signal length be $N$ samples and split it into $M$ segments at sample indices
@@
$h_{i,L}(\tau)=\mathrm{ISM}\bigl(\mathbf{x}^{s}_{i},\;\mathbf{x}^{r}_{i,L}\bigr),$ (7)
$h_{i,R}(\tau)=\mathrm{ISM}\bigl(\mathbf{x}^{s}_{i},\;\mathbf{x}^{r}_{i,R}\bigr).$ (8)
@@
Then the left and right binaural signals are obtained by block‐wise convolution:
$y_{L}[n]$ $=\sum_{i=0}^{M-1}\bigl(b_{i}*h_{i,L}\bigr)[\,n-n_{i}\,],$ (10)
$y_{R}[n]$ $=\sum_{i=0}^{M-1}\bigl(b_{i}*h_{i,R}\bigr)[\,n-n_{i}\,],$ (11)
where $h_{i,L}$ and $h_{i,R}$ are the left‐ and right‐ear impulse responses in $h_{i}$. Putting it all together, the final binaural signal is
@@
To ensure the output audio is in a valid range for playback, we normalize the signal to the range [1,1][-1,1]:
@@
By decomposing the task into three modular stages, our framework achieves high flexibility and extensibility. Each stage leverages either pretrained models or physical priors, allowing Sonic4D to naturally scale with future advances. This decoupled design further enables plug-and-play upgrades, making Sonic4D applicable to a wide range of immersive content creation scenarios.
Experiments
Experimental Settings
Since there is no existing benchmark for spatial audio in 4D scene exploration, we use two complementary evaluations.
SELD Evaluation.
Inspired by (Shimada et al. 2024), We adopt a pretrained Sound Event Localization and Detection(SELD) model (Diaz-Guerra et al. 2024), which extracts a Multi-ACCDOA representation from 5s of video+stereo audio (Shimada et al. 2022). Following (Diaz-Guerra et al. 2024), we evaluate on a curated subset of STARSS23 (Shimada et al. 2023), selecting perspective-view clips of musical instruments to avoid mosaic artifacts that may interfere with generative models.
We compare our method with the state-of-the-art ViSAGe model (Kim et al. 2025). As ViSAGe requires additional source location inputs (elevation $\theta$ and azimuth $\phi$), we use the global mean elevation from its training dataset (YT-Ambigen), and for azimuth, we utilize the ground-truth annotations provided in STARSS23. Following the approach in (Shimada et al. 2024), we convert the ViSAGe-generated FOA audio into stereo format by first rotating the audio according to the azimuth, and then applying a simple transformation: $left=W+Y$ and right=WYright=W-Y, where $W$ is the omnidirectional component and $Y$ is the first-order horizontal (left-right) component of the FOA audio.
To quantify spatial audio performance, we employ the joint localization and detection metrics from the DCASE SELD challenge (Diaz-Guerra et al. 2024; Shimada et al. 2024). Specifically, we report:
[…]
  • Directions of Arrival Angular Error ($\Delta_{\phi}$), the mean azimuth angular error. (Due to the SELD model, we only measure azimuth here.)
  • Distance Error ($\Delta_{d}$) and Relative Distance Error ($\Delta_{rd}$), measuring absolute and normalized distance estimation errors.
These metrics together capture both detection and localization quality in our spatial audio evaluations.
User Study.
@@
We choose MMAudio (Cheng et al. 2024) as the baseline method, whose generated audio is monaural and contains no spatialization. To determine whether our generated spatial audio conveys a sense of space, we paired the same viewpoint‑specific video with either the spatialized stereo audio or the non‑spatialized mono audio for comparison. Participants were first asked, for each video pair (identical visuals but different audio), to choose which audio track felt more spatial. Next, they rated both audio tracks on two criteria from 1 to 5:
@@
  • Spatial Localization Accuracy: Measures how faithfully the interaural level and time‐difference cues recreate the true 3D source position. Participants rate from 1 (“no sense of source direction”) to 5 (“precise and stable perception of source location”).
@@
  • Audio–Visual Spatial Alignment: Assesses how consistently the evolving audio cues (e.g. changes in loudness or binaural disparity) track the on‐screen motion of the subject or camera. Participants rate from 1 (“audio movement conflicts with visuals”) to 5 (“audio shifts perfectly match visual motion”).